# Long-Context Speculative Decoding with Efficient Drafting and Verification

### v1

@@ Abstract
Speculative decoding has become a promising technique to mitigate the high inference latency of autoregressive decoding in Large Language Models (LLMs). Despite its promise, the effective application of speculative decoding in LLMs still confronts three key challenges: the increasing memory demands of the draft model, the distribution shift between the short-training corpora and long-context inference, and inefficiencies in attention implementation. In this work, we enhance the performance of speculative decoding in long-context settings by addressing these challenges. First, we propose a memory-efficient draft model with a constant-sized Key-Value (KV) cache. Second, we introduce novel position indices for short-training data, enabling seamless adaptation from short-context training to long-context inference. Finally, we present an innovative attention aggregation method that combines fast implementations for prefix computation with standard attention for tree mask handling, effectively resolving the latency and memory inefficiencies of tree decoding. Our approach achieves strong results on various long-context tasks, including repository-level code completion, long-context summarization, and o1-like long reasoning tasks, demonstrating significant improvements in latency reduction. The code is available at https://github.com/sail-sg/LongSpec.
@@
[…]
Equal contribution
@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable success across various natural language processing tasks (Achiam et al., 2023), but their autoregressive decoding mechanism often results in high latency. To address this limitation, speculative decoding (Leviathan et al., 2023) has emerged as a promising solution. By employing a lightweight draft model to generate multiple candidate tokens, the target model can verify these tokens in parallel, thereby accelerating the inference process without compromising output quality.
Despite the great advancements in speculative decoding, existing research has primarily concentrated on short-context scenarios. However, as highlighted by Chen et al. (2025), the core potential of speculative decoding lies in long-context settings, where the maximum inference batch size is relatively low. The limited batch size restricts autoregressive decoding from fully utilizing the GPU computation resource, making speculative decoding an ideal approach to address this constraint. Yet, despite its advantages, the development of draft models specially designed for long-context scenarios remains largely unexplored.
While the need for a long-context draft model is clear from both application demands and theoretical considerations, we find that existing methodologies developed for short-context scenarios are inadequate when applied to longer sequences. This inadequacy stems from three emergent challenges unique to long-context speculative decoding: 1) Architecture: the extra memory overhead of the draft model, 2) Training: the distribution shift of position indices, and 3) Inference: the inefficiencies in tree attention implementation. First, as decoding length increases, previous State-of-The-Art (SoTA) autoregressive draft models like EAGLE (Li et al., 2024) and GliDe (Du et al., 2024) require linearly increasing KV caches, resulting in substantial storage overhead. This issue becomes critical in long-context settings, where memory usage is of vital importance. Second, the training data of the draft model mainly consists of short context data, rendering it undertrained over the large position indices (An et al., 2025), while the inference is for long-context data. The discrepancy will cause the draft model unable to perform speculation when facing large position indices of long-context input. Moreover, SoTA approaches often rely on tree attention, which is incompatible with the current advanced attention kernels due to the tree mask. This incompatibility further constrains the usage of tree speculative decoding in long-context scenarios.
To address these challenges, we introduce LongSpec, a framework for long-context speculative decoding, featuring architectural innovation (Sec. 3.1), novel training methods (Sec. 3.2), and optimized inference implementation (Sec. 3.3). The three key innovations significantly enhance the efficiency and scalability of speculative decoding in long-context scenarios.
First, to alleviate the memory overhead problem, we develop a draft model architecture that circumvents the linear expansion of KV caches which maintains a constant memory footprint as the context grows. Concretely, our approach employs a sliding window self-attention component to capture local dependencies, complemented by a cache-free cross-attention module for effectively modeling long-context representations. This approach effectively mitigates memory overhead which is particularly critical in long-context inference without compromising performance.
Next, to handle the training discrepancy problem, we propose the Anchor-Offset Indices to reconcile the short-context training for long-context inference. To fully train all the indices, we randomly add an offset to position indices. This ensures that some larger indices are also sufficiently trained. However, since RoPE (Su et al., 2024) is based on relative positions, directly adding an offset to all indices does not have a direct effect. Inspired by Streaming LLM (Xiao et al., 2024), we set the first four attention sink tokens as the anchor indices and only add the random offset to the remaining token index as shown in Figure 2. This indexing strategy ensures that all the indices can be sufficiently trained for the draft model. In contrast, the vanilla indexing strategy repeatedly trains only the smaller indices and is unable to train those exceeding the training set length. Meanwhile, anchoring the sink tokens ensures that the target model can approximate the attention sink patterns found in long texts, even with short texts.
Finally, to implement highly efficient tree attention, we propose a new computation method called Hybrid Tree Attention. Our insight comes from the discovery that the tree mask in tree-based speculative decoding can be decomposed into two parts, the previously cached part with a chain structure and the speculation part with a tree structure. Specifically, the tree mask is only required between the current queries and the speculation tokens (i.e., current input tokens) to ensure correctness. So we use Flash_Decoding (Dao, 2024) to compute the first part efficiently and use a custom Triton kernel fused_mask_attn to compute the second part. We then combine these components using a log-sum-exp trick, enabling our approach to accelerate tree attention computations up to 4.1 $\times$ compared to previous implementations in Medusa (Cai et al., 2024).
Extensive experiments are conducted to evaluate the effectiveness of LongSpec. Experiments on five long-context understanding datasets using five LLMs as target models show that our LongSpec can effectively reduce the long-context inference latency, leading to a maximum speedup of 3.26$\times$ compared with the strong baseline model with Flash_Decoding. Additional experiments on the long reasoning task AIME24 with the o1-like model QwQ (Qwen, 2024) further validate the effectiveness of LongSpec, achieving a 2.25$\times$ speedup in wall-clock time.
@@ 2 Related Work
Speculative decoding offers a promising approach to accelerating LLMs while maintaining the quality of their outputs. Early efforts, such as Speculative Decoding (Xia et al., 2023), SpS (Leviathan et al., 2023), BiLD (Kim et al., 2024), and OSD (Liu et al., 2024c), rely on existing smaller LLMs to generate draft sequences. Some other methods aim to improve upon those early efforts (Sun et al., 2023; Miao et al., 2024; Chen et al., 2024). There are also some works using part of the target model as the draft model (Liu et al., 2024a; Zhang et al., 2024; Elhoushi et al., 2024). Retrieval-based speculative decoding methods offer an alternative by utilizing $N$-gram matching rather than relying on smaller models. Examples include Lookahead Decoding (Fu et al., 2024), REST (He et al., 2024), and Ouroboros (Zhao et al., 2024). These approaches bypass the need for additional model training, leveraging pre-existing data patterns to construct draft sequences efficiently.
More recent advancements, including Medusa (Cai et al., 2024), EAGLE (Li et al., 2024), and GliDe (Du et al., 2024), have expanded on these foundations by designing specialized draft models and introducing tree-based speculative techniques. These methods leverage customized draft models tailored for speculative decoding, achieving higher efficiency and performance. Additionally, the tree-based approaches employed in these methods allow for more adaptive and parallelizable decoding processes, paving the way for broader applications in real-world systems.
Although speculative decoding has progressed significantly for conventional context lengths, only two existing papers focus on speculative decoding in long-context scenarios. TriForce (Sun et al., 2024) introduces a three-layer speculative decoding system that is scalable for long sequence generation. MagicDec (Chen et al., 2025) uses speculative decoding to improve both the throughput and latency of LLM inference. However, these methods mainly utilize the target model with the sparse KV cache as the draft model. The computation-intensive draft models restrict the practical usage of these methods when facing various batch sizes. In contrast, our work focuses on efficiently building a draft model with only one transformer block, achieving more effective performance across different scenarios.
@@ 3 Methodology
In this section, we present our framework LongSpec for Long-Context Speculative Decoding, which addresses three key challenges: (1) designing a lightweight draft model architecture with minimal additional memory overhead, (2) devising the training strategy with anchor-offset indices to handle long contexts effectively, and (3) implementing a fast tree attention mechanism that leverages tree-based speculation for practical usage. We detail each core component in the following subsections: Section 3.1 introduces our Memory-Efficient Architecture, Section 3.2 explains the Effective Training Regimes, and Section 3.3 describes the Fast Tree Attention implementation for inference.
@@ 3.1 Memory-Efficient Architecture
In previous work, the success of the SoTA model EAGLE depends on two factors: (1) the hidden states provided by the target model, and (2) an autoregressive structure. However, an autoregressive draft model inevitably needs to store its own KV cache, which introduces additional overhead in long-context inference requiring large GPU memory.
To avoid this extra memory overhead, we propose a draft model with constant memory usage regardless of the length of the context. As illustrated in Figure 1, our model comprises two components: the self-attention module and the following cross-attention module. The self-attention module focuses on modeling local context, while the cross-attention module captures long-context information. Because the self-attention module only processes local information, we adopt a sliding-window attention mechanism. Hence, during inference, the self-attention memory footprint does not exceed the window size, which we set to 512. We also provide the theoretical upper bound on the performance degradation caused by the slicing window in Section 3.4.
For the cross-attention component, inspired by GliDe (Du et al., 2024), we leverage the KV cache of the target model. This design not only enables better modeling of previous information but also completely removes additional storage overhead for long contexts, since the large model’s KV cache must be stored regardless of whether or not speculative decoding is employed. Different from GliDe, we also share the weights of the Embedding Layer and LM Head between the target model and the draft model, which significantly reduces the memory usage for large-vocabulary LLMs such as LLaMA.
Figure 1: Illustration of the memory-efficient draft model. We use sliding window self-attention to capture the local context information and cross attention layer to gather long-context information.
Figure 2: Illustration of the Anchor-Offset Indices and the Hybrid Tree Attention. (a) The left section illustrates the differences between the vanilla indexing and the Anchor-Offset Indices. By introducing a randomly selected offset and some anchor indices, the Anchor-Offset Indices enable the short-context training stage to seamlessly integrate with the long-context training stage. (b) The right section displays the Hybrid Tree Attention, which combines the advantages of Flash_Decoding and our Triton-implemented Attention.
@@ 3.2 Effective Training Regimes
Anchor-Offset Indices. With vanilla position indices, which consist of successive integers starting from 00, those indices appearing earlier in sequences occur more frequently than larger position indices (An et al., 2025), as shown in the Figure 2 upper left part. Consequently, larger position indices receive insufficient training updates, which leads to a training inference discrepancy. A common method to solve this problem is RoPE-based extrapolation, which trains the model with a smaller RoPE base and extends the RoPE base with interpolation for longer contexts (Gao et al., 2024; Liu et al., 2024d; Peng et al., 2024). However, directly using these methods will cause an inconsistent problem in draft model training. To leverage the target model’s KV cache, our draft model must keep the RoPE base the same as the target model. Based on our exploratory experiments, the inconsistent RoPE base between the target model and the draft model will cause a significant collapse of the cross-attention layer, which makes the draft model’s long-text capability degrade. The consistency requirement of RoPE base in the draft model limits the usage of methods like RoPE-based extrapolation, which requires flexible adjustments to the RoPE base.
Instead, we tackle this challenge by leveraging carefully designed indices. These indices must ensure that (1) the position indices in the draft model can be sufficiently trained using short-context data and (2) the indices would not cause the target model to exhibit out-of-distribution behavior because the target model shares the same indices as the draft model during training.
To satisfy these constraints, we propose the Anchor-Offset Indices strategy. Specifically, we reserve the first four positions $[0,1,2,3]$ as attention sink tokens (Xiao et al., 2024), then assign all subsequent tokens to large consecutive indices starting at a random offset (e.g., $[0,1,2,3,8192,8193,8194,\dots]$). The anchor indices and random offset ensure that every position index can be sufficiently trained, addressing the limitation of the vanilla one that repeatedly trains only smaller indices.
@@
Additionally, according to Xiao et al. (2024), LLM exhibits an attention sink phenomenon when dealing with long texts, which means the attention weights primarily concentrate on the first four tokens and the recent tokens. Therefore, we believe that utilizing Anchor-Offset Indices can naturally lead the target model to exhibit in-distribution behavior. In our experiments, adopting these indices in the target model only increases the loss by approximately 0.001, indicating that the target model is indeed well-suited to such changes. We also provide the theoretical upper bound of the distribution shift error in Section 3.4.
@@
Flash Noisy Training. During training our draft model leverages the KV caches from a large model, while these KV caches are not always visible during inference. This is because the large model only updates its KV cache upon verification completion. Concretely, for the cross-attention query $Q_{t}$ in the draft model, we can only guarantee access to the corresponding key-value states $K_{<t^{\prime}}$, $V_{<t^{\prime}}$ satisfying $1\leq|t^{\prime}-t|<\gamma$, where $\gamma$ is the number of speculative steps.
@@
To ensure consistency between training and inference, a straightforward solution would be to add an attention mask (Du et al., 2024). However, this method is incompatible with Flash Attention (Dao et al., 2023), which would significantly degrade training speed and cause prohibitive memory overhead, particularly in long-context training scenarios. Therefore, we propose a technique called flash noisy training. During training, we randomly shift the indices of queries and key-value states with $1\leq j<\gamma$. Suppose the sequence length is $l$, then we compute
$O_{\geq j}\;=\;\mathrm{flash\_attn}\bigl{(}Q_{\geq j},\,K_{<l-j},\,V_{<l-j}\bigr{)}.$

### v2

# LongSpec: Long-Context Lossless Speculative Decoding with Efficient Drafting and Verification

@@ Abstract
As Large Language Models (LLMs) can now process extremely long contexts, efficient inference over these extended inputs has become increasingly important, especially for emerging applications like LLM agents that highly depend on this capability. Speculative decoding (SD) offers a promising lossless acceleration technique compared to lossy alternatives such as quantization and model cascades. However, most state-of-the-art SD methods are trained on short texts (typically fewer than 4k tokens), making them unsuitable for long-context scenarios. Specifically, adapting these methods to long contexts presents three key challenges: (1) the excessive memory demands posed by draft models due to large Key-Value (KV) cache; (2) performance degradation resulting from the mismatch between short-context training and long-context inference; and (3) inefficiencies in tree attention mechanisms when managing long token sequences. This work introduces LongSpec, a framework that addresses these challenges through three core innovations: a memory-efficient draft model with a constant-sized KV cache; novel position indices that mitigate the training–inference mismatch; and an attention aggregation strategy that combines fast prefix computation with standard tree attention to enable efficient decoding. Experimental results confirm the effectiveness of LongSpec, achieving up to a 3.26$\times$ speedup over strong Flash Attention baselines across five long-context understanding datasets, as well as a 2.25$\times$ reduction in wall-clock time on the AIME24 long reasoning task with the QwQ model, demonstrating significant latency improvements for long-context applications.
@@
Figure 1: The SoTA SD method, EAGLE, has a training context length of 2048, which is significantly shorter than the context lengths of modern LLMs.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities (Achiam et al., 2023), and their ability to handle extensive contexts is becoming crucial for emerging applications such as LLM agents and long reasoning tasks (Tan et al., 2025; Guo et al., 2025), which now operate over context windows extending to millions of tokens (Team et al., 2024). In these demanding long-context scenarios, the high inference latency of standard autoregressive decoding becomes a pronounced bottleneck. While various acceleration techniques such as quantization (Lin et al., 2024), sparse attention (Li et al., 2024a), and model cascades (Gupta et al., 2024) have been proposed to mitigate this problem, they often compromise the output quality, rendering them lossy solutions. In contrast, speculative decoding (SD) (Leviathan et al., 2023) offers a lossless acceleration strategy by using a smaller draft model to propose token sequences, which are then verified in parallel by the larger target model. However, state-of-the-art (SoTA) SD methods (Li et al., 2024b), which often rely on a small and standalone draft model, are mainly designed and evaluated on short-context data, typically with sequences shorter than 4k tokens (see Figure 1). Although some existing SD methods can be extended to longer contexts, they often use the full target model with a compressed Key-Value (KV) cache as the draft model (Chen et al., 2025b; Tiwari et al., 2025). These approaches avoid the overhead of training a dedicated draft model, but their reliance on full target models, which are not sufficiently lightweight, limits the speed of draft generation. As a result, these methods may underperform compared to SoTA short-context SD techniques. This divergence raises a critical question:
Why cannot SoTA SD methods for short contexts be directly applied to long sequences?
In response to this question, we attribute the difficulty of directly adapting effective short-context SoTA SD techniques to long-context settings to three emergent challenges:
  1. 1.
    Architecture: In SoTA SD methods (e.g., EAGLE (Li et al., 2024b)), the draft model’s KV cache still grows linearly with context length. This linear growth becomes a prohibitive memory bottleneck as the context length increases.
  2. 2.
    Training: Language model training typically relies on plentiful short-sequence data, while long-sequence data remains relatively scarce. The imbalance of training data makes it difficult for the model to generalize to longer contexts. To address this, conventional wisdom in training long-context LLMs employs length extrapolation, in particular by extending the Rotary Position Embedding (RoPE) (Su et al., 2024) base to accommodate longer contexts (Gao et al., 2024; Liu et al., 2024d; Peng et al., 2024). However, this solution is not directly applicable to SoTA SD draft models, because their RoPE base must match that of the target model [^1] SoTA SD techniques often require the draft model to utilize intermediate features (e.g., hidden states or KV cache) from the target model, which is crucial for providing richer information from the target model, enabling the draft model to better align with and predict the target model’s outputs. See more explanations in Appendix B., which is fixed and already scaled for long-context scenarios.
  3. 3.
    Inference: The effectiveness of tree attention verification (Miao et al., 2024; Cai et al., 2024) diminishes in long-context scenarios. In particular, common inference optimizations for long-context scenarios are primarily designed to handle regular, structured attention masks and are not optimized for arbitrary or unstructured attention masks. As a result, potential speedups from speculation may be lower than expected.
To address these challenges, we introduce LongSpec, a comprehensive framework for efficient long-context lossless speculative decoding. LongSpec overcomes the aforementioned obstacles through three key innovations:
  1. 1.
    Memory-Efficient Architecture. We propose a draft model architecture with constant memory usage regardless of context length, effectively resolving the scalability limitations of prior SoTA autoregressive draft models.
  2. 2.
    Effective Training Regimes. We develop a novel training strategy involving Anchor-Offset Indices, enabling draft models trained on short sequences to robustly generalize to much longer contexts at inference time.
  3. 3.
    Fast Tree Attention. We introduce Hybrid Tree Attention, a new computation method that significantly speeds up tree verification by decomposing attention calculations and leveraging optimized Triton kernels.
Experiments on five long-context understanding datasets using five LLMs as target models show that our LongSpec can significantly reduce the long-context inference latency, achieving up to a 3.26$\times$ speedup over strong baselines with Flash Attention [^2] In this paper, Flash Attention refers to the inference optimization technique FlashDecoding (Dao et al., 2023), implemented via the flash_attn_with_kvcache function from the Flash Attention library (Dao, 2022)., and up to a 7$\times$ speedup over common baselines using the HuggingFace implementation. Additional experiments on the dataset AIME24 with the long reasoning model QwQ (Qwen, 2024) further validate the effectiveness of LongSpec, yielding a 2.25$\times$ speedup in wall-clock time. Furthermore, our proposed Anchor-Offset Indices enable models to reach the same loss level 3.93$\times$ faster, and our Hybrid Tree Attention reduces attention computation latency by approximately 75% compared to the standard HuggingFace implementation. The code is available at https://github.com/sail-sg/LongSpec
@@ 2 Related Work
Speculative decoding offers a promising approach to accelerating LLMs without compromising the quality of their outputs [^3] While this paper focuses on original speculative decoding methods which are lossless, some recent works explore lossy speculative decoding (see Appendix A for a brief overview).. Early efforts (Xia et al., 2023; Leviathan et al., 2023; Liu et al., 2024c; Bae et al., 2023) rely on existing smaller LLMs to generate draft sequences. Some other methods aim to improve upon those early efforts (Sun et al., 2023; Miao et al., 2024; Chen et al., 2024). There are also some works using part of the target model as the draft model (Liu et al., 2024a; Zhang et al., 2024; Elhoushi et al., 2024; Xia et al., 2025). Retrieval-based speculative decoding methods (Fu et al., 2024; He et al., 2024; Zhao et al., 2024) offer an alternative by utilizing $N$-gram matching rather than relying on smaller models. These approaches bypass the need for additional model training, leveraging pre-existing data patterns to construct draft sequences efficiently.
More recent advancements, including Medusa (Cai et al., 2024), EAGLE (Li et al., 2024b), and GliDe (Du et al., 2024), have expanded on these foundations by designing specialized draft models and introducing tree speculation and verification techniques. These methods leverage customized draft models tailored for speculative decoding, achieving higher efficiency and performance. Additionally, the tree-based approaches employed in these methods allow for more adaptive and parallelizable decoding processes, paving the way for broader applications in real-world systems.
Although speculative decoding has progressed significantly for conventional context lengths, only a few existing papers focus on lossless speculative decoding in long-context scenarios. TriForce (Sun et al., 2024) introduces a three-layer speculative decoding system that is scalable for long sequence generation. MagicDec (Chen et al., 2025b) uses speculative decoding to improve both the throughput and latency of LLM inference. QuantSpec(Tiwari et al., 2025) employs a hierarchical 4-bit quantized KV cache and 4-bit quantized weights for draft models. However, these methods mainly utilize the target model with the sparse KV cache as the draft model. The computation-intensive draft models restrict the practical usage of these methods when facing various batch sizes. In contrast, our work focuses on efficiently building a draft model with only one transformer block, achieving more effective performance across different scenarios.
Figure 2: Illustration of the memory-efficient draft model, the Anchor-Offset Indices, and the Hybrid Tree Attention. (a) We use a sliding window self-attention layer to capture the local context information and a cross-attention layer to gather long-context information. (b) The differences between the vanilla indexing and the Anchor-Offset Indices. By introducing a randomly selected offset and some anchor indices, the Anchor-Offset Indices enable the short-context training stage to seamlessly integrate with the long-context training stage. (c) The Hybrid Tree Attention combines the advantages of Flash Attention and our Triton-implemented Attention.
@@ 3 Methodology
In this section, we present our framework LongSpec for Long-Context Speculative Decoding, which addresses three key challenges by (1) designing a lightweight draft model architecture with constant-sized memory overhead, (2) devising the training strategy with anchor-offset indices to handle long contexts effectively, and (3) implementing a fast attention aggregation mechanism that leverages tree-based speculation and verification for practical usage.
@@ 3.1 Memory-Efficient Architecture
In previous work, the success of the SoTA model EAGLE depends on two key factors: (1) the hidden states provided by the target model, and (2) its autoregressive structure. However, an autoregressive draft model inevitably requires maintaining its own KV cache, which introduces additional overhead during long-context inference and demands substantial GPU memory, especially for tasks such as LLM agents and long reasoning that involve producing large amounts of output.
To avoid this extra memory overhead, we propose a draft model with constant memory usage independent of context length. As illustrated in Figure 2(a), our model consists of two components: the self-attention module and the following cross-attention module. The self-attention module focuses on modeling local context, while the cross-attention module captures long-range dependencies. To restrict memory usage, we apply a sliding-window attention mechanism to the self-attention module, a technique widely adopted in modern LLMs (Beltagy et al., 2020). Hence, during inference, the self-attention memory footprint does not exceed the window size, which we set to 512. In Section 3.4, we provide a theoretical upper bound on the performance degradation introduced by sliding-window attention compared to full-context self-attention. This bound ensures that the approximation maintains acceptable performance.
For the cross-attention component, inspired by GliDe (Du et al., 2024), we leverage the KV cache of the target model (see Appendix B for a detailed explanation of how this benefits the draft model). This design not only enables better modeling of previous information but also completely removes additional storage overhead for long contexts, since the large model’s KV cache must be stored regardless of whether or not speculative decoding is employed. Different from GliDe, we additionally share the weights of the Embedding Layer and LM Head between the target and draft models, which substantially reduces memory consumption for large-vocabulary LLMs such as LLaMA-3 (vocabulary size: 128,256) (Dubey et al., 2024) and Qwen-2.5 (vocabulary size: 152,064) (Yang et al., 2024).
@@ 3.2 Effective Training Regimes
Anchor-Offset Indices. With vanilla position indices, which consist of successive integers starting from 00, those indices appearing earlier in sequences occur more frequently than larger position indices (An et al., 2025), as shown in Figure 2(b) upper part. Consequently, larger position indices receive insufficient training updates, which leads to a training inference discrepancy. As we point out in Section 1, the common method, RoPE-based extrapolation, cannot be directly used here because the RoPE base is fixed when the target model is chosen and fixed. To leverage the target model’s KV cache, our draft model must keep the RoPE base the same as the target model. To tackle this challenge, we can only leverage carefully designed indices. These indices must ensure that (1) the position indices in the draft model can be sufficiently trained using short-context data and (2) the indices would not cause the target model to exhibit out-of-distribution behavior because the target model shares the same indices as the draft model during training.
@@
To satisfy these constraints, we propose the Anchor-Offset Indices strategy. Specifically, we reserve the first four positions $[0,1,2,3]$ as attention sink tokens [^4] According to (Xiao et al., 2024), LLM exhibits an attention sink phenomenon when dealing with long texts, which means the attention weights primarily concentrate on the first four tokens and the recent tokens., then assign all subsequent tokens to large consecutive indices starting at a random offset (e.g., $[0,1,2,3,8192,8193,8194,\dots]$). By exploiting the attention sink phenomenon, we believe that utilizing Anchor-Offset Indices can naturally lead the target model to exhibit in-distribution behavior. The anchor indices and random offset ensure that every position index can be sufficiently trained, addressing the limitation of the vanilla one that repeatedly trains only smaller indices. In our experiments, adopting these indices in the target model only increases the loss by approximately 0.001, indicating that the target model is indeed well-suited to such changes. We further present a theoretical upper bound for the learning error of the draft model in Section 3.4, which accounts for a position distribution shift term. Notably, the expression of this term suggests that our method incurs fewer errors compared to the vanilla position indices.
@@
Flash Noisy Training. During training, our draft model leverages the KV cache from a large model, while this KV cache is not always visible during inference. This is because the large model only updates its KV cache upon verification completion. Concretely, for the $t$-th cross-attention query $Q_{t}$ in the draft model, we can only guarantee access to the corresponding key-value states $K_{<t^{\prime}}$, $V_{<t^{\prime}}$ satisfying $1\leq|t^{\prime}-t|<\gamma$, where $\gamma$ is the number of speculative steps.
@@
To ensure consistency between training and inference, a straightforward solution would be to add an attention mask (Du et al., 2024). However, this method is incompatible with Flash Attention, which would significantly degrade training speed and cause prohibitive memory overhead, particularly in long-context training scenarios. Therefore, we propose a technique called flash noisy training. During training, we randomly shift the indices of queries and key-value states with $1\leq j<\gamma$. Suppose the sequence length is $l$, then we compute
$O_{\geq j}=\mathsf{attn}\bigl{(}Q_{\geq j},\,K_{<l-j},\,V_{<l-j}\bigr{)}.$